a projectile is fired from a point at125m/s so as to strike a point a horizontal distance of 100m and 200m higher than the point of firing neglecting air resistance, calculate (a) the angle with the horizontal, at which the projectile should be fired in order to strike the point in minimum time (b) the time taken by the projectile to strike the point


1
Expert's answer
2022-02-15T16:14:16-0500

"y=-\frac{gx^2}{2v^2\cos^2\alpha}+x\tan\alpha,"

"x=100,~y=200,"

"\alpha=65.4°,"

"t=\frac{x}{v\cos\alpha}=1.9~s."